OpenStudy (dehelloo):

What are the y and x intercepts for -2x^2 + 7x -3 For the y intercept I got (0, -3) But for the x intercepts using the general quad formula I could not get an answer as I cant square a negative. Not sure what I'm doing wrong - could someone give it a go please.

OpenStudy (gorv):

y=-2x^2 + 7x -3 for x intercept put y=0 -2x^2 + 7x -3 =0 -2x^2 + 6x+x -3=0 (x-3)*(-2x+1)=0 x=3 or 1/2 y intercept put x=0 y=-3

OpenStudy (dehelloo):

Thanks for the reply but I'm not sure what you are doing here. How do you get from -2x^2 + 7x -3 =0 to -2x^2 + 6x+x -3=0

OpenStudy (anonymous):

@dehelloo x=3&1/3 y=-3. wet get this by putting x=0(for y-intercept& y=0(for x-intercept).

OpenStudy (dehelloo):

Yep I understand that we have to turn x to = for the y intercept and y into 0 for the x int. I am just not sure what to do after that.

OpenStudy (campbell_st):

he has split the middle term into 2 parts that can be factored... -2x^2 + 6x + x - 3 = 0 factor each pair or terms -2x^2 + 6x has common factors of -2x so you get -2x(x - 3) + 1(x -3) = 0 now you have a common factor of (x -3) so you end up with (x -3)(-2x + 1) = 0 so you now have the binomial factors of the quadratic... just solve each binomial... for the x intercepts hope this makes sense

OpenStudy (campbell_st):

so you now need to find the values of x that make either binomial equal to zero so you are solving x - 3 = 0 and -2x + 1 = 0 so when y = 0 x will equal your solutions

OpenStudy (dehelloo):

How does he split the middle term into 2 parts.

OpenStudy (dehelloo):

Where does he get the extra x from and how does the 7 change into a 6.

OpenStudy (anonymous):

hi guys